-- HarryRenshall - 08 Feb 2008

Week of 080211

Open Actions from last week:

Understand why afs31 failure stopped cms LSF staging jobs

Monday:

See the weekly joint operations meeting

Tuesday:

Present in room: Harry, Roberto, Simone, Jamie, Flavia, Gavin. James Present on phone: Derek, Michael, Damielel, Gonzalo, Stephen

elog review:

  • FTS:
    • Proxy issue witrh FTS. 2 distimcty issues, not yet understood
    • A couple of pther FTS software issues (not yet understood) causing inefficiencies, stuck transfers.
    • Proper solution will require a software fix. Workaround pending.

  • IN2P3 mkdir issue on dCache repoted to dCache developers (ticket number 2509).
    • Will require software fix.

Experiment report(s):

ATLAS:

  • FDR data done OK.
  • CCRC'08 data starting tomorrow: T0->T1 export according to MoU shares. Target 200Hz for 10 hours ~ 500-600 MB/s average export from CERN (peak at 1.1 GB/s)
  • Plan to schedule simultaneous run with CMS later this week

CMS:

  • Week 2 plan: ramp up Castor based sites for T0->T1.
  • T1-> T1 concentrate on CNAF, RAL, ASGC exporting to other T1s.

LHCb:

  • Pit -> T0: working well, T0->tape: working well
  • T0 -> T1: 350MB/s working well.
  • T1 re-processing: expt. software is not ready yet.

Alice:

  • migration problems, xrootd-disk/tape filesize mismatch problems to be understood

Core services (CERN) report: nothing more to report.

  • Q. High load noted in morning daily meeting: any issues seen from this? No.

DB services (CERN) report:

  • The streams propagation to PIC stopped last Friday 08.02 for 20 minutes, due to a disk failure of the ATLAS and LHCb databases at PIC and re- synchronized when the services restarted at PIC.

Monitoring / dashboard report: nothing new to report.

Release update: nothing.

Questions/comments from sites/experiments:

  • Gonzalo (PIC): LHcb - which dashboard should we use? Roberto: Gridview for now. DIRAC3 is missing a dashboard.

AOB: nothing.

Wednesday

elog review:

Experiment report(s):

Core services (CERN) report:

DB services (CERN) report:

Monitoring / dashboard report:

Release update:

Questions/comments from sites/experiments:

AOB:

Thursday

elog review:

Experiment report(s):

Core services (CERN) report:

DB services (CERN) report:

Monitoring / dashboard report:

Release update:

Questions/comments from sites/experiments:

AOB:

Friday

elog review:

Experiment report(s):

Core services (CERN) report:

DB services (CERN) report:

Monitoring / dashboard report:

Release update:

Questions/comments from sites/experiments:

AOB:

Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2008-02-13 - GavinMcCance
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback